# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v2

@@ 4.5 Performance Evaluation in Reduced-Complexity Environments: MiniGrid
While the CREATE environment (Jain et al.,, 2020) showcased SETLE’s ability to learn from rich, dynamic, and continuous visual environments, a central hypothesis of our work is that the same underlying mechanism can generalise across varying representational regimes—including low-resolution, symbolic environments. To test this, we extended SETLE to the MiniGrid benchmark suite, a grid-based environment that offers highly abstract, discrete observations and minimal visual complexity.
@@
To mitigate the limitations of low-resolution symbolic input in MiniGrid, further refinements were applied to the object extraction pipeline. Rendered frames were first pre-processed using contrast equalisation and spatial upscaling to enhance visual saliency. Each segmented crop was then passed through CLIP (Ramesh et al.,, 2022) to obtain a robust vision-language embedding. CLIP was chosen explicitly over conventional encoders such as ResNet (He et al.,, 2016) or VGG (Simonyan and Zisserman,, 2014) because these standard architectures, trained on natural image statistics, catastrophically fail when applied to synthetic grid-based environments with flat colours and uniform textures. CLIP, by contrast, demonstrated the ability to align symbolic patterns with semantic concepts even in low-visual-complexity domains. This design choice was critical for preserving the object-centric nature of the SETLE framework and maintaining consistency between the original CREATE experiments and the MiniGrid adaptation.
@@ 4.6 Statistical Analysis and Error Reporting
To rigorously evaluate our results, we performed a statistical analysis on the final epoch losses from 8 successful training runs. The data did not significantly deviate from a normal distribution, as confirmed by a Shapiro-Wilk test (p = 0.156), justifying the use of parametric tests. We grouped the runs by their margin value ($\alpha$) into three methodologically sound, discrete categories: Low ($\alpha\leq 0.5$), Medium ($0.5\leq\alpha\leq 1.0$), and High ($\alpha\geq 1.0$).
A one-way ANOVA was first conducted to compare the effect of the margin group on performance. While the omnibus test did not reveal a statistically significant difference between the three groups (F(2, 5) = 2.34, p = 0.178), a large effect size was observed (partial $\eta^{2}$ = = 0.48), suggesting a strong underlying trend, and that statistical significance may have been obscured due to a limited sample size and the arbitrary assignment of values to each category: low, medium, or high.
To investigate this trend further, we conducted a planned comparison between the extreme groups. An independent samples t-test confirmed that the ’Low’ margin group produced a statistically significant lower mean loss (M=0.71, SD=0.05) compared to the ’High’ margin group (M=0.85, SD=0.55), t(5) = -2.83, p = 0.037. This result provides strong evidence that lower margin values lead to more stable and effective training for the encoder in our experiments.
@@ 5.1 Overview of the Baseline RL Architecture
The baseline agent is based on a Double Deep Q-Network (Double DQN) (Van Hasselt et al.,, 2016), selected for its stability in Q-value estimation and compatibility with mixed observation and action modalities. The agent uses two Q-heads: one for discrete action selection and one for continuous control, depending on the environment.
@@ 5.3 Training Regimes and Experimental Design
Each configuration is tested on five CREATE tasks and a suite of MiniGrid scenarios. Experiments are conducted across three random seeds per condition to ensure statistical robustness. All runs are logged using Weights & Biases (W&B), capturing reward trajectories, Q-value trends, enrichment frequency, and memory match statistics.
@@ 5.3.1 Results in Physically Grounded Tasks (CREATE)
Out of over 40 independent runs across different tasks and configurations, only those using both the adapter and soft update enhancements managed to solve the environment at least once (Fig. 26). These results highlight not only the potential of structured trajectory enrichment, but also the fragility of the approach in complex domains. Importantly, these experiments were conducted under deliberately constrained training regimes: episode lengths were capped at 9 steps, and overall training time was limited. We deliberately constrained the training setup—limiting the number of episodes and capping each trajectory at 9 steps—to focus on evaluating whether structured memory and enrichment mechanisms can accelerate learning and enable generalisation under low-data, few-step conditions. This design tests the efficiency of SETLE-enhanced agents in challenging settings where standard reinforcement learning strategies would typically struggle to succeed without extensive trial-and-error or curriculum-based shaping.
Figure 26: The agent’s performance was tested across more than 40 independent runs. The chart visualizes a key finding from these experiments: across all 40+ runs, the Baseline, Action Selection Only, and SETLE Optimisation strategies never succeeded once. Only the ’Adapter and Penalty’ strategy achieved a non-zero success rate.
@@ 5.3.2 Results in Symbolic Reasoning Tasks (MiniGrid)
Rather than hallucinating reward where none exists, SETLE retrieves structured prior experiences—such as object co-occurrences and contextual affordances, that allow the agent to infer when sequences are not worth propagating optimistic returns. In doing so, the agent suppresses speculative Q-value growth and aligns its internal value function with the actual utility of its behaviours. This form of value realism is increasingly recognised as a desirable property in deep RL. Recent work (Fujimoto et al.,, 2018; Kumar et al.,, 2020) has highlighted the risks of overestimation in Q-learning and proposed conservative or uncertainty-aware alternatives to promote trustworthy generalisation.
@@ 6 Discussion
A key challenge in RL is to ensure that learnt representations transfer across tasks rather than remaining task-specific (D’Eramo et al.,, 2024; Chen et al.,, 2024). By incorporating hierarchical trajectory representations, we enable the recognition of shared patterns and support knowledge transfer. This aligns with findings that hierarchical structuring is an essential mechanism for generalisation (Sancho et al.,, 2024) and reflects principles from human cognition, where learning depends on recombining meaningful substructures from experience (Lake et al.,, 2017). Many AI models, however, struggle with compositional flexibility, limiting their adaptability. The importance of SETLE’s hierarchical structure was empirically validated; our ablation studies demonstrated that flattening the graph representation substantially degraded clustering quality, with the Dunn Index for success clusters dropping from 1.652 to 1.326 (Table 3). This shows that our framework advances a more effective, hierarchical approach that enables agents to retain and reuse structured knowledge.
@@
Although hierarchical structures have been central to RL research for a long time (Botvinick et al.,, 2009), most current approaches depend on predefined skill segmentation, which limits the ability to autonomously discover reusable structures. In contrast, by modelling the relational dependencies between trajectory components, our method aligns with broader trends in structured representation learning (Hafner et al.,, 2019; Roads and Love,, 2024), where explicitly modelling these relationships enhances transferability.
@@
In addition to generalising tasks, this work also advances lifelong learning and adaptive AI. Effective open-ended learning depends on a structured memory system that enables agents to accumulate and refine knowledge over time (Finn et al.,, 2017). The hierarchical memory structure in SETLE lays the groundwork for such capabilities. This adaptability extends across domains; while object-level embeddings failed to align between the CREATE and MiniGrid environments, SETLE’s trajectory-level representations revealed functional similarities, such as a mean cosine similarity of 0.34 between CreateLevelBelt-v0 and MiniGrid-Empty-5x5-v0 (Fig. 35), demonstrating that the framework learns transferable task abstractions. This structured approach also mitigates the issue of catastrophic forgetting.
@@
This work contributes to the broader discussion on structured intelligence in AI (Mohan et al.,, 2024; Eckstein and Collins,, 2020) by providing an effective mechanism to link low-level state-action execution with high-level conceptual reasoning. Achieving human-like adaptability requires moving beyond mere pattern recognition to extract, manipulate, and reuse structured representations across various domains (Schneider and Logan,, 2006; Hommel et al.,, 2000; Wu et al.,, 2022). While similar approaches in RL have been proposed, challenges still exist in designing frameworks that can generalise beyond task-specific representations (Mohan et al.,, 2024).
@@
While our framework offers several advantages, a number of challenges remain. One key limitation is scalability; future work should evaluate SETLE in more complex, real-world environments such as robotics simulations, which involve richer dynamics and higher-dimensional state representations. This, however, poses significant challenges, as the computational complexity of encoding such graphs can be prohibitive, and the efficiency of hierarchical representations in these tasks is an unresolved issue. Additionally, this work primarily focuses on representation learning. A key area for future research is to deepen the understanding of how these structured representations impact policy optimisation. Moving forward, exploring how SETLE can be extended toward dynamic adaptation, enabling agents to continuously refine their representations in response to evolving task demands, will be crucial for pushing reinforcement learning toward lifelong, generalisable intelligence.

### v3

@@ 4.5 Performance Evaluation in Reduced-Complexity Environments: MiniGrid
While the CREATE environment (jain2020generalization) showcased SETLE’s ability to learn from rich, dynamic, and continuous visual environments, a central hypothesis of our work is that the same underlying mechanism can generalise across varying representational regimes—including low-resolution, symbolic environments. To test this, we extended SETLE to the MiniGrid benchmark suite, a grid-based environment that offers highly abstract, discrete observations and minimal visual complexity.
@@
To mitigate the limitations of low-resolution symbolic input in MiniGrid, further refinements were applied to the object extraction pipeline. Rendered frames were first pre-processed using contrast equalisation and spatial upscaling to enhance visual saliency. Each segmented crop was then passed through CLIP (ramesh2022hierarchical) to obtain a robust vision-language embedding. CLIP was chosen explicitly over conventional encoders such as ResNet (he2016deep) or VGG (simonyan2014very) because these standard architectures, trained on natural image statistics, catastrophically fail when applied to synthetic grid-based environments with flat colours and uniform textures. CLIP, by contrast, demonstrated the ability to align symbolic patterns with semantic concepts even in low-visual-complexity domains. This design choice was critical for preserving the object-centric nature of the SETLE framework and maintaining consistency between the original CREATE experiments and the MiniGrid adaptation.
@@ 4.6 Statistical Analysis and Error Reporting
To rigorously evaluate our results, we performed a statistical analysis on the final epoch losses from 8 successful training runs. The data did not significantly deviate from a normal distribution, as confirmed by a Shapiro-Wilk test (N=8, p = 0.156), justifying the use of parametric tests. We grouped the runs by their margin value ($\alpha$) into three methodologically sound, discrete categories: Low ($\alpha\leq 0.5$), Medium ($0.5\leq\alpha\leq 1.0$), and High ($\alpha\geq 1.0$).
A one-way ANOVA was conducted to compare the effect of margin group on performance (N=8). The omnibus test did not reach conventional statistical significance (F(2, 5) = 2.34, p = 0.178). This p-value corresponds to approximately 82% evidence against the null hypothesis, indicating moderate but sub-threshold support for group differences. Given the limited sample size (8 runs across 3 groups), this analysis is likely underpowered. A large partial $\eta^{2}$ (0.48) was observed and is reported descriptively; however, no inferential conclusions are drawn from effect size in the absence of statistical significance. These results are therefore exploratory and motivate further experiments with increased repetitions and formal power analysis.
To investigate this trend further, we conducted a planned comparison between the extreme groups. An independent samples t-test confirmed that the ’Low’ margin group (N=3) produced a statistically significant lower mean loss (M=0.71, SD=0.05) compared to the ’High’ margin group (N=3, M=0.85, SD=0.55), t(5) = -2.83, p = 0.037. This result provides strong evidence that lower margin values lead to more stable and effective training for the encoder in our experiments.
@@ 5.1 Overview of the Baseline RL Architecture
The baseline agent is based on a Double Deep Q-Network (Double DQN) (van2016deep), selected for its stability in Q-value estimation and compatibility with mixed observation and action modalities. The agent uses two Q-heads: one for discrete action selection and one for continuous control, depending on the environment.
@@ 5.3 Training Regimes and Experimental Design
Each configuration is tested on five CREATE tasks and a suite of MiniGrid scenarios. Experiments are conducted across three independent runs with different PyTorch random seeds to ensure statistical robustness. All runs are logged using Weights & Biases (W&B), capturing reward trajectories, Q-value trends, enrichment frequency, and memory match statistics.
Randomness was handled using the default pseudo-random number generators provided by each library. Python’s built-in random module and NumPy’s global RNG rely on Mersenne Twister (MT19937), while PyTorch uses a Philox-based generator for CUDA operations. PyTorch seeds were explicitly set for each run, while Python and NumPy relied on their default generators. Each training run was executed independently, and no seed hashing or reuse was applied. We acknowledge that the number of repetitions is limited and that future work will incorporate explicit seeding across all libraries and increased repetitions to enable more rigorous statistical analysis.
@@ 5.3.1 Results in Physically Grounded Tasks (CREATE)
Out of 42 independent runs across different tasks and configurations, only those using both the adapter and soft update enhancements managed to solve the environment at least once (Fig. 26). These results highlight not only the potential of structured trajectory enrichment, but also the fragility of the approach in complex domains. Importantly, these experiments were conducted under deliberately constrained training regimes: episode lengths were capped at 9 steps, and overall training time was limited. We deliberately constrained the training setup—limiting the number of episodes and capping each trajectory at 9 steps—to focus on evaluating whether structured memory and enrichment mechanisms can accelerate learning and enable generalisation under low-data, few-step conditions. This design tests the efficiency of SETLE-enhanced agents in challenging settings where standard reinforcement learning strategies would typically struggle to succeed without extensive trial-and-error or curriculum-based shaping.
Figure 26: The agent’s performance was tested across 42 independent runs. The chart visualizes a key finding from these experiments: across all 42 runs, the Baseline, Action Selection Only, and SETLE Optimisation strategies never succeeded once. Only the ’Adapter and Penalty’ strategy achieved a non-zero success rate.
@@ 5.3.2 Results in Symbolic Reasoning Tasks (MiniGrid)
Rather than hallucinating reward where none exists, SETLE retrieves structured prior experiences—such as object co-occurrences and contextual affordances, that allow the agent to infer when sequences are not worth propagating optimistic returns. In doing so, the agent suppresses speculative Q-value growth and aligns its internal value function with the actual utility of its behaviours. This form of value realism is increasingly recognised as a desirable property in deep RL. Recent work (fujimoto2018addressing; kumar2020conservative) has highlighted the risks of overestimation in Q-learning and proposed conservative or uncertainty-aware alternatives to promote trustworthy generalisation.
@@ 6 Discussion
A key challenge in RL is to ensure that learnt representations transfer across tasks rather than remaining task-specific (deramo2024; chen2024). By incorporating hierarchical trajectory representations, we enable the recognition of shared patterns and support knowledge transfer. This aligns with findings that hierarchical structuring is an essential mechanism for generalisation (sancho2024) and reflects principles from human cognition, where learning depends on recombining meaningful substructures from experience (lake2017building). Many AI models, however, struggle with compositional flexibility, limiting their adaptability. The importance of SETLE’s hierarchical structure was empirically validated; our ablation studies demonstrated that flattening the graph representation substantially degraded clustering quality, with the Dunn Index for success clusters dropping from 1.652 to 1.326 (Table 3). This shows that our framework advances a more effective, hierarchical approach that enables agents to retain and reuse structured knowledge.
@@
Although hierarchical structures have been central to RL research for a long time (botvinick2009hierarchically), most current approaches depend on predefined skill segmentation, which limits the ability to autonomously discover reusable structures. In contrast, by modelling the relational dependencies between trajectory components, our method aligns with broader trends in structured representation learning (hafner2020dream; roads2024dimensions), where explicitly modelling these relationships enhances transferability.
@@
In addition to generalising tasks, this work also advances lifelong learning and adaptive AI. Effective open-ended learning depends on a structured memory system that enables agents to accumulate and refine knowledge over time (finn2017model). The hierarchical memory structure in SETLE lays the groundwork for such capabilities. This adaptability extends across domains; while object-level embeddings failed to align between the CREATE and MiniGrid environments, SETLE’s trajectory-level representations revealed functional similarities, such as a mean cosine similarity of 0.34 between CreateLevelBelt-v0 and MiniGrid-Empty-5x5-v0 (Fig. 35), demonstrating that the framework learns transferable task abstractions. This structured approach also mitigates the issue of catastrophic forgetting.
@@
This work contributes to the broader discussion on structured intelligence in AI (mohan2024structuredRL; eckstein2020hierarchicalRL) by providing an effective mechanism to link low-level state-action execution with high-level conceptual reasoning. Achieving human-like adaptability requires moving beyond mere pattern recognition to extract, manipulate, and reuse structured representations across various domains (schneider2006; hommel2000; wu2022). While similar approaches in RL have been proposed, challenges still exist in designing frameworks that can generalise beyond task-specific representations (mohan2024structuredRL).
While our framework offers several advantages, several challenges remain. Although we have adapted the HeCo architecture as the foundation for SETLE’s encoder, we have not incorporated modified versions of heterogeneous GNN baselines (e.g., HeCo, GTC) as direct comparison methods in our experimental evaluation. As discussed in Section 4.1, this decision was based on fundamental differences in problem formulation. Whereas SETLE addresses episodic, dynamic graphs representing trajectories, methods like HeCo and GTC are designed for static, node-centric graph tasks. This limitation reflects a broader challenge in the field: the existing divide between static graph mining and dynamic reinforcement learning. Most state-of-the-art heterogeneous GNNs, including HeCo (wang2021selfsupervised) and the more recent GTC (sun2025gtc), are optimised for node-centric tasks on invariant graphs. Adapting these architectures to serve as standalone baselines for episodic, subgraph-centric trajectory encoding remains an open challenge. Without standardised evaluation protocols that bridge static representation learning and sequential decision-making, it remains difficult to quantify how much an agent’s performance stems from the underlying graph architecture versus the way it structures its experience over time. While this limitation does not detract from the validity of the present work, which demonstrates SETLE’s effectiveness in encoding structured trajectories and improving RL performance, developing such unified benchmarks is a necessary step toward understanding how structured representations can be decoupled from specific task dynamics to achieve more generalizable intelligence
@@
Another key limitation is scalability; future work should evaluate SETLE in more complex, real-world environments such as robotics simulations, which involve richer dynamics and higher-dimensional state representations. This, however, poses significant challenges, as the computational complexity of encoding such graphs can be prohibitive, and the efficiency of hierarchical representations in these tasks is an unresolved issue. Additionally, this work primarily focuses on representation learning. A key area for future research is to deepen the understanding of how these structured representations impact policy optimisation. Moving forward, exploring how SETLE can be extended toward dynamic adaptation, enabling agents to continuously refine their representations in response to evolving task demands, will be crucial for pushing reinforcement learning toward lifelong, generalisable intelligence.